OpenStudy (anonymous):

Solve the system algebraically. 2x+ y - 10 = 0 x - y - 4 = 0 What is the value of y?

OpenStudy (anonymous):

Okay, because this is a system, it's not just two equations. y will be the same for both equations. 2x + y - 10 = 0 x - y - 4 = 0 You want to solve for y, so you have to multiply the bottom equation by -2 so that you can eliminate x when you add them together. 2x + y - 10 = -2x + 2y + 8 y - 10 = 2y + 8 -y - 10 = 8 -y = 18 y = -18

OpenStudy (anonymous):

Thanks ikram.